### v5

@@ 4.2 Robustness to alternative game forms
We now look at specifications with different numbers of states of the world, different utility specifications and different distributions over the states of the world.
In Figure 11 we consider simulations with $n=3$ and $n=9$ states of the world, so that any two adjacent states are spaced $0.5$ and $0.125$ from each other, respectively. The figure confirms that all conclusions reached in the previous section extend to the case of a smaller or larger state space. However, with a large number of states, behaviour tends to be more noisy. This is explained by the relative increase in exploration due to the change in size of the agents’ Q-matrices. In general, as we keep $\lambda$ fixed to the base case configuration, each state-action pair is on average visited more (less) often depending on the size of the agent’s Q-matrix. This eventually results in improving (worsening) the agent’s learning.
@@
Finally, in Figure 12 we show outcomes for different distributions over the states of the world; namely, a probability distribution with linearly increasing probability mass, and one with linearly decreasing probability mass. In this case results also indicate communication is roughly in line with the most optimistic theoretical benchmark. Now, however, existence of the ex-ante optimal equilibrium is no longer guaranteed. Hence, we rely on the receiver-preferred partitional equilibrium as comparator. A surprising result is obtained in the case of the decreasing distribution. While in all our simulations agents do better than babbling, here the receiver obtains a payoff lower than the babbling one, while the sender obtains a larger one. We think this finding is interesting because the sender seems able to manipulate the receiver even when theoretically it should not be possible and the receiver is losing out from not just playing the ex-ante optimal action. Unfortunately, we do not have a convincing explanation for this result. We believe it might be due to the existence of multiple optimal actions for the receiver even when the sender’s strategy is partitional.
Figure 10: Ex-ante expected reward for the sender (left) and receiver (right) for different levels of bias. Cases with $3$ states (top) and $9$ states (bottom).
@@
Figure 12: Ex-ante expected reward for the sender (left) and receiver (right) for different levels of bias. Cases with a linearly increasing distribution (top) and linearly decreasing distribution (bottom). We use $p(\theta_{k})$ to indicate the probability mass on the $k$-th state in $\Theta=\{\theta_{1},\ldots,\theta_{n}\}$. There are $n=6$ states as in the base-case simulations.

### v6

@@ 4.2 Robustness to alternative game forms
We now look at specifications with different numbers of types, different utility specifications and different distributions over the types.
In Figure 11 we consider simulations with $n=3$ and $n=9$ types, so that any two adjacent types are spaced $0.5$ and $0.125$ from each other, respectively. The figure confirms that all conclusions reached in the previous section extend to the case of a smaller or larger type space. However, with a large number of types, behaviour tends to be more noisy. This is explained by the relative decrease in exploration due to the change in size of the agents’ Q-matrices. In general, as we keep $\lambda$ fixed to the base case configuration, each state-action pair is on average visited more (less) often the smaller (larger) the agent’s Q-matrix is. This eventually results in improving (worsening) the agent’s learning.
@@
Finally, in Figure 12 we show outcomes for different distributions over the types; namely, a probability distribution with linearly increasing probability mass, and one with linearly decreasing probability mass. In this case results also indicate communication is roughly in line with the most optimistic theoretical benchmark. Now, however, existence of the ex-ante optimal equilibrium is no longer guaranteed. Hence, we rely on the receiver-preferred partitional equilibrium as comparator. A surprising result is obtained in the case of the decreasing distribution. While in all our simulations agents do better than babbling, here the receiver obtains a payoff lower than the babbling one, while the sender obtains a larger one. We think this finding is interesting because the sender seems able to manipulate the receiver even when theoretically it should not be possible and the receiver is losing out from not just playing the ex-ante optimal action. Unfortunately, we do not have a convincing explanation for this result. We believe it might be due to the existence of multiple optimal actions for the receiver even when the sender’s strategy is partitional.
Figure 10: Ex-ante expected reward for the sender (left) and receiver (right) for different levels of bias. Cases with $3$ types (top) and $9$ types (bottom).
@@
Figure 12: Ex-ante expected reward for the sender (left) and receiver (right) for different levels of bias. Cases with a linearly increasing distribution (top) and linearly decreasing distribution (bottom). We use $p(\theta_{k})$ to indicate the probability mass on the $k$-th type in $\Theta=\{\theta_{1},\ldots,\theta_{n}\}$. There are $n=6$ types as in the base-case simulations.